Find the ratio of: 
 30 days to 36 hours
Find the ratio of: 
 30 days to 36 hours

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We know that,
1 day = 24 hours
Then,
30 days = 30 × 24 = 720 hours
Now we have to find the ratio,
= 720/36
= 20/1... [∵divide both by 36] So, the required ratio is 20: 1.

Thanks